2. Solve the following exponential equation. Exact answers only.
49^x-12*7^x=-36

Answer by lwsshak3(11628)

2. Solve the following exponential equation. Exact answers only.
49^x-12*7^x=-36
7^2x-12*7^x+36=0
let u=7^x
u^2=7^2x
u^2-12u+36=0
(u-6)(u-6)=0
u=6=7^x
xlog7=log6
x=log6/log7